OpenStudy (anonymous):

find the derivative... f(x) = sqrt(x^2+5)

OpenStudy (anonymous):

\[f(x)=\sqrt(x^2+5)\]

OpenStudy (anonymous):

actually find the equation of the tangent line of the above function at [2,3]

sam (.sam.):

\[f(x)=(x^2+5)^{1/2}\] Power rule followed by chain rule

sam (.sam.):

What do you get?

sam (.sam.):

When you use power rule, you should get \[f(x)=(x^2+5)^{1/2} \\ \\ f'(x)=\frac{1}{2}(x^2+5)^{-1/2}[\frac{d}{dx}(x^2+5)]\] Where \(\frac{d}{dx}(x^2+5)\) is the result of the chain rule

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but I never learned about d/dx

sam (.sam.):

That is just telling you to differentiate x^2+5

OpenStudy (anonymous):

all I know is the f(x+h) - f(x) /h

sam (.sam.):

That's another method, but way confusing to solve this, you'll have x and h after differentiating

OpenStudy (anonymous):

so the derivative of \[f(x)=\sqrt(x^2-5)\] is just \[\frac{ 1 }{ 2 }(x^2-5)^{1/2}\]

OpenStudy (anonymous):

negative 1/2

sam (.sam.):

\[\frac{ 1 }{ 2 }(x^2\color{red}{+}5)^{-1/2}\]

sam (.sam.):

You still need to multiply by the differential of \(x^2+5\) according to chain rule, the final result is \[f'(x)=\frac{1}{2}(x^2+5)^{-1/2}[\frac{d}{dx}(x^2+5)] \\ \\ f'(x)=\frac{1}{2}(x^2+5)^{-1/2}(2x) \\ \\ f'(x)=x(x^2+5)^{-1/2} \]

OpenStudy (anonymous):

ok, i think i got it now, but with other method, I just multiply \[\sqrt{(x+h)^2-5}-\sqrt{x^2-5}\] by the conjugate pair \[\frac{ \sqrt{(x+h)^2-5}+\sqrt{x^2-5} }{\sqrt{(x+h)^2-5}+\sqrt{x^2-5}}\]

sam (.sam.):

Ok did you got it with your method?

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\]\[\frac{ 1 }{ \sqrt{x^2+5} }\]

OpenStudy (anonymous):

?

sam (.sam.):

\[f'(x)=\frac{ x }{ \sqrt{x^2+5} }\]

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

ok, equation of tangent line is \[y=3-\frac{ 1 }{ 6 }(x-2)\] at the point [2,3]

sam (.sam.):

Gotta go I'll be back later

OpenStudy (anonymous):

well im done thanks!